OpenStudy (anonymous):

Write the terms of the series and find their sum ∑k=14(k+2)2

OpenStudy (anonymous):

wait wrong equation

OpenStudy (goformit100):

@samtasticc

OpenStudy (anonymous):

\[\sum_{k=1}^{4} (k+2)\]

OpenStudy (anonymous):

\[(k+2)^{2}\]

OpenStudy (anonymous):

\[\sum_{k=1}^4(k+2)^2\]??

OpenStudy (anonymous):

put \(k=1\) get \((1+2)^2=3^2=9\) put \(k=2\) get \((2+2)^2=4^2=16\) etc then add up the four numbers

OpenStudy (anonymous):

Yes thats correct

OpenStudy (anonymous):

i.e. add up \(9+16+25+36\)

OpenStudy (anonymous):

so you keep going like k=3 and k=4 and then those four numbers added together is the answer?

OpenStudy (anonymous):

What are the terms of the series? 9,16,25 & 36 @satellite73